# Covariance-Based Device Activity Detection with Massive MIMO for Near-Field Correlated Channels

### v1

@@ Abstract
This paper studies the device activity detection problem in a massive multiple-input multiple-output (MIMO) system for near-field communications (NFC). In this system, active devices transmit their signature sequences to the base station (BS), which detects the active devices based on the received signal. In this paper, we model the near-field channels as correlated Rician fading channels and formulate the device activity detection problem as a maximum likelihood estimation (MLE) problem. Compared to the traditional uncorrelated channel model, the correlation of channels complicates both algorithm design and theoretical analysis of the MLE problem. On the algorithmic side, we propose two computationally efficient algorithms for solving the MLE problem: an exact coordinate descent (CD) algorithm and an inexact CD algorithm. The exact CD algorithm solves the one-dimensional optimization subproblem exactly using matrix eigenvalue decomposition and polynomial root-finding. By approximating the objective function appropriately, the inexact CD algorithm solves the one-dimensional optimization subproblem inexactly with lower complexity and more robust numerical performance. Additionally, we analyze the detection performance of the MLE problem under correlated channels by comparing it with the case of uncorrelated channels. The analysis shows that when the overall number of devices is large or the signature sequence length is small, the detection performance of MLE under correlated channels tends to be better than that under uncorrelated channels. Conversely, when is small or is large, MLE performs better under uncorrelated channels than under correlated ones. Simulation results demonstrate the computational efficiency of the proposed algorithms and verify the correctness of the analysis.
@@
In this paper, we study the covariance-based activity detection problem under the spatial correlated Rician fading channel, where the deterministic line-of-sight (LoS) components and correlation matrices are derived from near-field spherical wave propagation [14, 15, 16, 17]. This setup makes both algorithm design and detection performance analysis more challenging than those under traditional uncorrelated Rayleigh fading channels, primarily due to the channel correlation. For spatially uncorrelated channels, the MLE formulation relies on the marginal distributions of each column of the received signal, as these columns are independently distributed [6]. However, in the case of spatially correlated channels, different columns of the received signal are dependent with each other and it is necessary to consider their joint distribution. The joint distribution complicates the MLE problem because its dimension is much higher than that of the marginal distributions and increases with the number of antennas. Moreover, since the existing detection performance analyses [6, 7, 8] assume that different columns of the received signal are independent and identically distributed (i.i.d.), they cannot be directly applied to the case of spatially correlated channels.
@@ I-B Main Contributions
On the algorithmic side, we propose two computationally efficient algorithms for solving the MLE problem: an exact CD algorithm and an inexact CD algorithm. The exact CD algorithm iteratively solves the one-dimensional optimization subproblem exactly using matrix eigenvalue decomposition and polynomial root-finding. However, due to its high complexity and potential numerical errors when the rank of the channel correlation matrix is large, the exact CD algorithm is not efficient and sometimes is even divergent. To address this issue, we propose the inexact CD algorithm, which primarily relies on Taylor expansion to approximate some terms in the objective function. The resulting inexact CD algorithm has a lower complexity and is guaranteed to converge. Additionally, we study a special case where only a few devices’ channels are correlated with a low-rank correlation matrix, while the channels of many other devices are uncorrelated. We demonstrate that exploiting the low-rank properties of the correlation matrices significantly reduces the complexity of both the exact CD and inexact CD algorithms.
@@ I-C Paper Organization and Notations
The rest of the paper is organized as follows. Section II introduces the system model and the MLE formulation. Section III presents the efficient CD algorithms for solving the MLE problem. In Section IV, we analyze the difference in detection performance between correlated and uncorrelated channels. Simulation results are provided in Section V. Finally, we conclude the paper in Section VI.
@@ II-A System Model
We consider an uplink single-cell massive MIMO system consisting of a BS equipped with an -antenna uniform linear array with half-wavelength antenna spacing. In the cell, there are single-antenna devices, and only devices are active during any coherence interval due to the sporadic traffic. We use to denote the binary variable with for active and for inactive devices. We consider the case where the devices are in the near-field region, i.e., the distance between the devices and the BS is less than the Rayleigh distance, which is given by where and denote the antenna aperture and the wavelength, respectively [14]. We assume that both the deterministic LoS channel and statistical multipath channels exist, and the channel between device and the BS can be modeled as the following Rician fading model [14, Eq. (133)]:
where is the LoS channel between device and the BS, and
where denotes the channel gain between device and the BS, is the distance between the device and the -th antenna, and is the distance between the device and the midpoint of the antenna array. The correlation matrix is given by [14, Eq. (134)]:
where is the number of scatterers between device and the BS, is the intensity attenuation caused by the -th scatterer, denotes the channel gain between the -th scatterer and the device and is the channel vector between the -th scatterer and the BS with a similar structure to (2), i.e., where and denote the corresponding channel gain and distances. Notice that and thus is a low-rank matrix when the number of scatterers
Each device is preassigned a unique signature sequence with length for activity detection. During the uplink pilot stage, all active devices synchronously transmit their signature sequences to the BS as random access requests. Following this, the received signal at the BS can be expressed as
@@ II-B Problem Formulation
In this paper, we focus on the algorithm design and performance analysis for the MLE problem (10). Notice that the dimension of the covariance matrix in problem (10) is which is significantly different from the case of uncorrelated channels. In the latter case, the covariance matrix involved in the MLE formulation is [5, 6, 7]. This difference makes problem (10) much more challenging to solve.
III Proposed CD Algorithms for Device Activity Detection
In this section, we focus on designing efficient algorithms for solving the MLE problem (10). Notice that problem (10) is nonconvex due to the concavity of We propose two algorithms for solving the MLE problem: an exact CD and an inexact CD. In the exact CD algorithm, we solve the one-dimensional optimization subproblem exactly, while in the inexact CD algorithm, we solve it inexactly in order to reduce the computational cost. Furthermore, we study a special case that allows us to further reduce the complexity of both the exact CD and inexact CD algorithms by utilizing the low-rank property of the near-field channel correlation matrix.
@@ III-A1 Solving Problem (19) Exactly
where and are the derivatives of and respectively. Notice that the left-hand side in (24) is a polynomial of degree so problem (24) can be solved using a polynomial root-finding algorithm [56, Chap. 6]. Next, we choose the solution with the smallest objective function value in (23) among the roots of (24) and the two boundary points and This yields the exact solution to problem (23) (i.e., problem (19)). The procedure of solving problem (19) exactly is outlined in lines 6 and 7 of Algorithm 1.
The above proposed method is suitable to solve problem (19) for small due to the following reasons. When is large, significant numerical errors arise in two steps: (i) the eigenvalue decomposition of i.e., the computation of and satisfying (20), and (ii) the root-finding of the polynomial of degree in (24). The error in step (i) makes the coefficients of and inaccurate, while the error in step (ii) makes it impossible to obtain a solution to problem (23). To address these issues, we propose a method for solving problem (19) inexactly, as detailed below.
@@ III-A2 Solving Problem (19) Inexactly
Theorem 1 demonstrates that, at each coordinate update, although the inexact CD algorithm does not solve problem (19) exactly, it finds a good approximate solution that sufficiently decreases the objective function, especially when is large. This guarantees the convergence of the inexact CD algorithm (to a stationary point).
@@ III-A3 Complexity
In massive MIMO systems, updating and storing dominate the complexity of both the exact and inexact CD algorithms. As follows, we consider a special case where the complexity of updating and storing can be significantly reduced.
@@ III-B A Special Case Study
In this subsection, we consider a special case in which the proposed exact CD and inexact CD can be accelerated. The motivation arises from the fact that within a cell, there may be only a few devices in the near-field region of the BS, whereas the majority of devices are located in the far-field region. Furthermore, we assume that for each device in the near-field region, the number of scatterers is small. In this scenario, we can model the near-field channel as correlated channel with low-rank correlation matrix (since ), and assume that the far-field channel is uncorrelated. In this scenario, we only focus on reducing the complexity of updating and storing by taking advantage of the low-rank properties of the near-field correlation matrices. Below, we formally present the assumption for this specific case.
@@
Notice that Then, a sufficient condition for Assumption 1 to hold is that there are few near-field devices (i.e., is small) and few scatterers in the near-field channels (i.e., is small). The following proposition demonstrates that under Assumption 1, () can be transformed into a matrix where only the upper-left submatrix is nonzero.
@@ IV Detection Performance Analysis
For each device , the channel follows where is the large-scale fading coefficient.
@@ IV-A Signal Statistical Dimension
First we define the signal statistical dimensions and in Case I and Case II respectively:
@@ IV-B Identifiability of True Activity Vector
The phase transition analysis in [7] demonstrates that, a crucial condition for the MLE model (10) to perform effectively is that the true activity indicator vector needs to be uniquely identifiable. This implies that there should be no other vector such that We now investigate the identifiability of in both Case I and Case II. Since both and are multivariate Gaussian distributions, their equivalence implies that their corresponding means and covariances must be identical. Therefore, the identifiability of is equivalent to the condition that there is no other such that and Furthermore, we define
@@
where follows from the Cauchy-Schwartz inequality and equality holds if and only if there exists a constant such that We observe from (54) that and are closer to orthogonal than and implying that the mutual interference between devices and in Case I is lower than that in Case II. The above analysis demonstrates that the channel correlation among different antennas can reduce the correlation between devices and thereby improve the detection performance. Therefore, when the signal statistical dimensions in Case I and Case II are equal (i.e., is large), the MLE model (10) exhibits superior detection performance in Case I (than in Case II).
V Simulation Results
@@
We consider an mMTC system with a single cell of radius m, where all potential devices are uniformly distributed. In the simulation, the carrier frequency is set as GHz. The scatterers are uniformly distributed within m from the BS, and the intensity attenuation is for all The channel gain and follow the path-loss model, which is given by where is the corresponding distance in km. We adopt a power control scheme such that where the power is set as dBm. The background noise power is set as dBm/Hz over MHz. The signature sequences are generated randomly and uniformly from the discrete set as in [8].
V-A Computational Efficiency of Proposed Algorithms
@@
In this subsection, we showcase the efficiency of the proposed exact and inexact CD algorithms for solving the device activity detection problem (10). In the inexact CD algorithm, we set the parameter to
@@
In Fig. 1, we show the decrease of the objective function values of the proposed exact and inexact CD algorithms with respect to the running time with and The outputs of the exact and inexact CD algorithms are denoted as and respectively. We then define as the minimum of and We observe from Fig. 1(a) that when the objective function values of exact and inexact CD can gradually decrease and converge to the same value. Moreover, exact CD is significantly more efficient than inexact CD. This is because, when is small (implying is also small), the exact CD algorithm can find the exact solution to the one-dimensional optimization subproblem (19) in each coordinate update, whereas the inexact CD algorithm can only find an approximate solution. In Fig. 1(b), we observe that when the objective function value of the inexact CD algorithm still converges, but that of the exact CD algorithm cannot decrease and diverges. This occurs because when is large (corresponding to a large ), the exact CD algorithm fails to solve the one-dimensional subproblem (19) due to the numerical errors associated with the high-dimensional eigenvalue decomposition and polynomial root-finding. This observation shows that exact CD is more efficient when is small, while inexact CD is more robust when is large.
@@
In Fig. 2, we assume that Assumption 1 holds true and plot the decrease of the objective function values of the proposed exact and inexact CD algorithms as well as their accelerated versions by Algorithm 2 with respect to the running time with and In Fig. 2(a), we set in Assumption 1, and for (implying that ). Consequently, we have in (30). Notice that in this case, there are devices with channel correlation matrix Fig. 2(a) shows that, in this scenario, the objective function value of the accelerated inexact CD converges to and it is more computationally efficient than the inexact CD algorithm. However, neither exact CD nor its accelerated version converges. This is because neither exact CD nor its accelerated version can solve the one-dimensional optimization subproblem (19) when the channel correlation matrix of the device To demonstrate that the exact CD algorithm can also be accelerated when it converges, we remove devices with channel correlation matrix Therefore, in Fig. 2(b), we set in Assumption 1, and for all We observe from Fig. 2(b) that the accelerated version of exact CD converges, and it is more efficient than exact CD. The same observation applies to the inexact CD algorithm as well.
@@
A summary of the proposed CD algorithms is as follows. First, the exact CD algorithm is more computationally efficient than the inexact CD algorithm when the rank of the channel correlation matrix is small. Second, the inexact CD algorithm is more favorable in terms of robustness in the case where the rank of the channel correlation matrix is large. Finally, both of the exact and inexact CD algorithms can be accelerated when Assumption 1 holds. Due to its robustness in solving the subproblem, we adopt the inexact CD algorithm to solve problem (10) in different scenarios in the following two subsections.
V-B Numerical Validation of Detection Performance Analysis

### v2

@@ Abstract
This paper studies the device activity detection problem in a massive multiple-input multiple-output (MIMO) system for near-field communications (NFC). In this system, active devices transmit their signature sequences to the base station (BS), which detects the active devices based on the received signal. In this paper, we model the near-field channels as correlated Rician fading channels and formulate the device activity detection problem as a maximum likelihood estimation (MLE) problem. Compared to the traditional uncorrelated channel model, the correlation of channels complicates both algorithm design and theoretical analysis of the MLE problem. On the algorithmic side, we present the classical exact coordinate descent (CD) algorithm for solving the MLE problem, which suffers from numerical instability when applied to correlated channels. We propose a computationally efficient inexact CD algorithm by approximating the objective function, which approximately solves the one-dimensional subproblem and improves both computational efficiency and numerical stability. Additionally, we analyze the detection performance of the MLE problem under correlated channels by comparing it with the case of uncorrelated channels. The analysis shows that when the overall number of devices is large or the signature sequence length is small, the detection performance of MLE under correlated channels tends to be better than that under uncorrelated channels. Conversely, when is small or is large, MLE performs better under uncorrelated channels than under correlated ones. Finally, we study the MLE model in the joint device activity and data detection context. Simulation results demonstrate the computational performance of the presented algorithms and verify the correctness of the analysis.
@@
In this paper, we study the covariance-based activity detection problem under the spatially correlated Rician fading channel, where the deterministic line-of-sight (LoS) components and correlation matrices for the near-field devices are derived from near-field spherical wave propagation [15, 16, 17, 18]. This setup makes both algorithm design and detection performance analysis more challenging than those under traditional uncorrelated Rayleigh fading channels, primarily due to the channel correlation. For spatially uncorrelated channels, the MLE formulation relies on the marginal distributions of each column of the received signal, as these columns are independently distributed [7]. However, in the case of spatially correlated channels, different columns of the received signal are dependent with each other and it is necessary to consider their joint distribution. The joint distribution complicates the MLE problem because its dimension is which is much larger than that of the marginal distribution (i.e., ), where is the signature sequence length and is the number of antennas. This is particularly relevant in the context of massive MIMO. Moreover, since the existing detection performance analyses [7, 8, 9] assume that different columns of the received signal are independent and identically distributed (i.i.d.), they cannot be directly applied to the case of spatially correlated channels.
@@ I-B Main Contributions
On the algorithmic side, we introduce an exact CD algorithm for solving the MLE problem, which is a direct extension of the CD algorithm designed for the uncorrelated channel case [6]. It exactly solves the one-dimensional optimization subproblem by using matrix eigenvalue decomposition and polynomial root-finding techniques. However, this algorithm suffers from numerical instability and can fail to converge when the channel correlation matrix has a high rank. To address this issue, we propose an inexact CD algorithm, which primarily relies on Taylor expansion to approximate selected terms in the objective function. This approach provides improved numerical stability and reduced per-coordinate update complexity. Additionally, we study a special case where only a few devices’ channels are correlated with a low-rank correlation matrix, while the channels of many other devices are uncorrelated. We demonstrate that exploiting the low-rank properties of the correlation matrices significantly reduces the complexity of both the exact CD and inexact CD algorithms.
@@
In addition, we study the joint device activity and data detection problem, where each active device wishes to send a few bits of information to the BS. In this scenario, each device is associated with multiple signature sequences and can embed a small number of information bits. We propose the MLE model for this scenario. Simulation results show that transmitting several bits of information does not significantly deteriorate the detection performance of the MLE model, especially when the length of the signature sequence is large.
@@ I-C Paper Organization and Notations
The rest of the paper is organized as follows. Section II introduces the system model and the MLE formulation. Section III presents the efficient CD algorithms for solving the MLE problem. In Section IV, we analyze the difference in detection performance between correlated and uncorrelated channels. In Section V, we study the MLE model in the joint device activity and data detection context. Simulation results are provided in Section VI. Finally, we conclude the paper in Section VII.
@@ II-A System Model
We consider an uplink single-cell massive MIMO system consisting of a BS equipped with an -antenna uniform linear array with half-wavelength antenna spacing. In the cell, there are single-antenna devices, and only devices are active during any coherence interval due to the sporadic traffic. We use to denote the binary variable with for active and for inactive devices. We consider the case where some devices are located in the near-field region of the BS, and others in the far-field region, as shown in Fig. 1. The radius of the near-field region is indicated by the Rayleigh distance, which is given by where and denote the antenna aperture and the wavelength, respectively [15]. Notice when the carrier frequency is set as GHz, and the number of antennas is set as and the corresponding Rayleigh distances are approximately m, m, and m, respectively. Consequently, the number of near-field devices can be large due to the substantial Rayleigh distance.
We assume that both the deterministic LoS channel and statistical multipath channels exist, and the channel between device and the BS can be modeled as the following Rician fading model:
where is the LoS channel between device and the BS. For a near-field device the mean vector can be expressed as [15, Eq. (133)]:
where denotes the channel gain between device and the BS, is the distance between the device and the -th antenna, and is the distance between the device and the midpoint of the antenna array. Furthermore, the correlation matrix is given by [15, Eq. (134)]:
where is the number of scatterers between device and the BS, is the intensity attenuation caused by the -th scatterer, denotes the channel gain between the -th scatterer and the device and is the channel vector between the -th scatterer and the BS with a similar structure to (2), i.e., where and denote the corresponding channel gain and distances. Notice that and thus is a low-rank matrix when the number of scatterers For a far-field device the vector can be represented by the conventional plane wave model, and the correlation matrix can be modeled as where is the large-scale fading coefficient [44].
Each device is preassigned a unique signature sequence with length for activity detection. During the uplink pilot stage, all active devices synchronously transmit their signature sequences to the BS as random access requests. Following this, the received signal at the BS can be expressed as [6, 7, 8]
@@ II-B Problem Formulation
When solving the MLE problem (10) for device activity detection, we only need to know the mean and the correlation matrix of each device’s channel, without requiring the knowledge of the channel vector i.e., the channel state. Notice that appears in (4) because of the system model [6, 7, 8], and appears in (10) due to the probability distribution function of the complex Gaussian distribution [57].
In this paper, we focus on the algorithm design and performance analysis for the MLE problem (10). Notice that the dimension of the covariance matrix in problem (10) is which is much larger than that in the case of uncorrelated channels [6, 7, 8], i.e., Therefore, the complexity of updating and storing the matrix in the CD algorithm is much higher than that in the case of uncorrelated channels. In addition, in the case of uncorrelated channels, updating a single coordinate of corresponds to a rank- update of the covariance matrix, which enables a closed-form solution for each coordinate update [6, 7, 8]. However, in the correlated case, due to (II-B), the rank- update property does not hold, making it impossible to obtain a closed-form solution for coordinate updates. These differences make problem (10) much more challenging to solve.
III Efficient CD Algorithms for Device Activity Detection
In this section, we focus on designing efficient algorithms for solving the MLE problem (10). Notice that problem (10) is nonconvex due to the concavity of We first present the exact CD algorithm for solving the MLE problem, which solves the one-dimensional optimization subproblem exactly. In contrast, the proposed inexact CD algorithm solves this subproblem approximately, thereby achieving improved numerical stability and reduced computational cost. Furthermore, we study a special case that allows us to further reduce the complexity of both the exact CD and inexact CD algorithms by utilizing the low-rank property of the near-field channel correlation matrix.
@@ III-A1 Solving Problem (19) Exactly
where and are the derivatives of and respectively. Notice that the left-hand side in (24) is a polynomial of degree so problem (24) can be solved using a polynomial root-finding algorithm [58, Chap. 6]. Next, we choose the solution with the smallest objective function value in (23) among the roots of (24) and the two boundary points and This yields the exact solution to problem (23) (i.e., problem (19)). The procedure of solving problem (19) exactly is outlined in lines 6 and 7 of Algorithm 1.
The above method is only suitable to solve problem (19) for small As observed in Section VI-A, the exact CD algorithm may diverge when is large. The observed divergence can be attributed to two main factors. First, for large the structures of and may cause their values to become very close to zero during certain coordinate updates. As a result, even small numerical errors in computing and can lead to significant inaccuracies in the objective function value of problem (23). Second, the root-finding process in problem (24) can produce numerically inaccurate solutions when the leading coefficient of the polynomial becomes excessively large, a scenario that becomes increasingly likely as increases. To address these issues, we propose a method for solving problem (19) inexactly, as detailed below.
@@ III-A2 Solving Problem (19) Inexactly
Theorem 1 demonstrates that, at each coordinate update, although the inexact CD algorithm does not solve problem (19) exactly, it finds a good approximate solution that sufficiently decreases the objective function, especially when is large. Furthermore, building upon the techniques introduced in [59], this theorem also establishes the convergence of the inexact CD algorithm to a stationary point.
@@ III-A3 Complexity
In massive MIMO systems, updating and storing dominate the complexity of both the exact and inexact CD algorithms. In the next subsection, we consider a special case where the complexity of updating and storing can be significantly reduced.
III-A4 Differences between Presented Algorithms and Those in [9]
First, we emphasize that the optimization problem (10) is significantly different from problem (7) in [9]. This difference is primarily due to two key factors: (i) the dimension of the covariance matrix is significantly larger than that in [9]; and (ii) the rank-1 update property, which holds for the covariance matrices in [9], does not apply to
Next, we compare the exact CD algorithm with the CD algorithm in [9]. Both of them transform the optimization problem into a polynomial root-finding problem, but unlike [9] where coefficients are computed analytically, this paper proposes applying eigenvalue decomposition to determine the polynomial coefficients.
Finally, we compare the inexact CD algorithm with the inexact CD algorithm in [9]. The main difference between the two algorithms is in their handling of the parameter In [9], the parameter is dynamically adjusted at each coordinate update. However, in this paper, the parameter is fixed to a sufficiently large positive number to ensure convergence. This difference arises because in [9], the computational complexity of the objective function is implying that adjusting does not incur any additional computational cost. In contrast, in this paper, adjusting would result in a computational complexity for the objective function of at least Consequently, we fix to reduce the complexity of coordinate updates.
@@ III-B A Special Case Study
In this subsection, we consider a special case in which the presented exact CD and inexact CD can be accelerated. The motivation arises from the fact that within a cell, there may be only a few devices in the near-field region of the BS, whereas the majority of devices are located in the far-field region. Furthermore, we assume that for each device in the near-field region, the number of scatterers is small. In this scenario, we can model the near-field channel as correlated channel with low-rank correlation matrix (since ), and assume that the far-field channel is uncorrelated. In this scenario, we only focus on reducing the complexity of updating and storing by taking advantage of the low-rank properties of the near-field correlation matrices. Below, we formally present the assumption for this specific case.
@@
Notice that Then, the limitation to Assumption 1 is twofold: first, the number of devices whose channels are correlated needs to be sufficiently smaller than the number of antennas; second, their channel correlation matrices must be low-rank. For example, consider a scenario where the number of antennas at the BS is There are potential devices in the cell, but only devices are in the near-field region, and the channels of the other devices are uncorrelated. Assuming the number of scatterers in the near-field channel is (i.e., ), we then have
The following proposition demonstrates that under Assumption 1, () can be transformed into a matrix where only the upper-left submatrix is nonzero.
@@ IV Detection Performance Analysis
For each device the channel follows where is the large-scale fading coefficient.
@@ IV-A Signal Statistical Dimension
The main idea of this subsection is to define and compare signal statistical dimensions in Case I and Case II. We infer that a larger signal statistical dimension corresponds to better detection performance because the received signal can be transformed by a unitary transformation. In the transformed signal vector, only the covariance matrix of a subvector provides information about The length of this subvector is the signal statistical dimension. We then compare the upper bounds of the signal statistical dimensions in Case I and Case II, which indicates the relationship between the system parameters and the signal statistical dimensions.
We now define the signal statistical dimensions and in Case I and Case II respectively:
@@ IV-B Identifiability of True Activity Vector
The main idea of this subsection arises from the phase transition analysis in [8], which demonstrates that a crucial condition for the MLE model (10) to perform effectively is that the true activity indicator vector needs to be uniquely identifiable. This implies that there should be no other vector such that We demonstrate that the unique identifiability of is more likely to be achieved in Case I than in Case II. Furthermore, inspired by the unique identifiability of we infer that the mutual interference between different devices is weaker in Case I compared to Case II.
We now investigate the identifiability of in both Case I and Case II. Since both and are multivariate Gaussian distributions, their equivalence implies that their corresponding means and covariances must be identical. Therefore, the identifiability of is equivalent to the condition that there is no other such that and Furthermore, we define
@@
where follows from the Cauchy-Schwartz inequality and equality holds if and only if there exists a constant such that We observe from (54) that and are closer to orthogonal than and implying that the mutual interference between devices and in Case I is lower than that in Case II. The above analysis demonstrates that the channel correlation among different antennas can reduce the mutual interference between devices and thereby improve the detection performance. Therefore, when the signal statistical dimensions in Case I and Case II are equal (i.e., is large), the MLE model (10) exhibits superior detection performance in Case I (than in Case II).
V Joint Device Activity and Data Detection
This section aims to show that the MLE model (10) can be applied to the joint device activity and data detection problem, in which each device is associated with multiple signature sequences and can embed a few information bits [10].
Suppose that each active device intends to send bits of information to the BS, where is a small integer. To perform joint activity and data detection, we assign each device a unique signature sequence set of size :
where When device is active and needs to send bits of data, it selects a single sequence from its assigned set and sends this sequence to the BS. The BS then detects which sequences have been transmitted to perform the device activity detection and data decoding.
The process of performing joint activity and data detection using the covariance-based approach is as follows. Let indicate whether sequence of device (i.e., ) is transmitted. Since at most one sequence is selected by each device, satisfies where indicates that device is active, and indicates that device is inactive. Then, the received signal at the BS, as well as its vectorization, can be expressed as follows:
Therefore, we have where the mean is
and the covariance matrix is given by
Using the same techniques presented in Section II-B, the MLE problem can be expressed as
To make the problem more computationally tractable, we relax the binary constraint to as in [7, 8, 9], and we relax the combination constraint as in [10]. Then we obtain the following relaxed problem:
Therefore, by solving problem (60) using the exact CD and inexact CD algorithms presented in this paper, we can effectively perform joint activity and data detection.
VI Simulation Results
@@
We consider an mMTC system with a single cell of radius m, where all potential devices are uniformly distributed. In the simulation, the carrier frequency is set as GHz. The scatterers are uniformly distributed within m from the BS, and the intensity attenuation is for all The channel gain and follow the path-loss model, which is given by where is the corresponding distance in km. We adopt a power control scheme such that where the power is set as dBm. The background noise power is set as dBm/Hz over MHz. The signature sequences are generated randomly and uniformly from the discrete set as in [9]. Notice that there are enough devices in the near-field region because, as mentioned in Section II-A, the Rayleigh distance can be sufficiently large, especially when the number of antennas is large.
All experiments are implemented in Matlab 2021b on a computer equipped with an Intel Core i7-11700 CPU and 32 GB of memory. This is for a fair comparison in computational complexity with existing algorithms. In practical implementations, more efficient programming languages such as C/C++ would be adopted, and the code would also be optimized for even more efficient execution.
VI-A Computational Performance of Presented Algorithms
@@
In this subsection, we demonstrate the computational performance of the exact and inexact CD algorithms for solving the device activity detection problem (10). In the inexact CD algorithm, we set the parameter to
Table I shows the convergence behavior of the exact CD algorithm for different ’s. For each we record the proportion of experiments in which the exact CD algorithm converges out of experiments. We observe that for the system parameter settings listed in Table I, the exact CD algorithm always converges when (corresponding to ). Conversely, the algorithm always diverges when (for most cases ). Comparing Setting 1 with Setting 2, we observe that as the number of antennas decreases from to the convergence threshold for increases from to Similarly, comparing Setting 1 and Setting 3, we find that reducing the signature sequence length from to also increases the convergence threshold for from to Furthermore, by comparing Setting 1 with Settings 4 and 5, we observe that parameters and do not influence the convergence threshold of
@@
In Fig. 2, we show the decrease of the objective function values of the exact and inexact CD algorithms with respect to the running time with and The outputs of the exact and inexact CD algorithms are denoted as and respectively. We then define as the minimum of and We observe from Fig. 2 that when and the objective function values of exact and inexact CD can gradually decrease and converge to the same value. Moreover, exact CD is significantly more efficient than inexact CD. This is because, when is small (implying is also small), the exact CD algorithm can find the exact solution to the one-dimensional optimization subproblem (19) in each coordinate update, whereas the inexact CD algorithm can only find an approximate solution. In Fig. 2(d), we observe that when the objective function value of the inexact CD algorithm still converges, but that of the exact CD algorithm cannot decrease and diverges. This observation shows that exact CD is more efficient when is small, while inexact CD is more robust when is large.
@@
In Fig. 3, we assume that Assumption 1 holds true and plot the decrease of the objective function values of the exact and inexact CD algorithms as well as their accelerated versions by Algorithm 2 with respect to the running time with and In Fig. 3(a), we set in Assumption 1, and for (implying that ). Consequently, we have in (30). Notice that in this case, there are devices with channel correlation matrix Fig. 3(a) shows that, in this scenario, the objective function value of the accelerated inexact CD converges to and it is more computationally efficient than the inexact CD algorithm. However, neither exact CD nor its accelerated version converges. This is because neither exact CD nor its accelerated version can solve the one-dimensional optimization subproblem (19) when the channel correlation matrix of the device To demonstrate that the exact CD algorithm can also be accelerated when it converges, we remove devices with channel correlation matrix Therefore, in Fig. 3(b), we set in Assumption 1, and for all We observe from Fig. 3(b) that the accelerated version of exact CD converges, and it is more efficient than exact CD. The same observation applies to the inexact CD algorithm as well.
@@
A summary of the presented CD algorithms is as follows. First, the exact CD algorithm is more computationally efficient than the inexact CD algorithm when the rank of the channel correlation matrix is small. Second, the inexact CD algorithm is more favorable in terms of robustness in the case where the rank of the channel correlation matrix is large. Finally, both of the exact and inexact CD algorithms can be accelerated when Assumption 1 holds. Due to its robustness in solving the subproblem, we adopt the inexact CD algorithm to solve problem (10) in different scenarios in the following two subsections.
VI-B Numerical Validation of Detection Performance Analysis